# Learning to Rank Critical Road Segments via Heterogeneous Graphs with OD Flow Integration

### v2

@@ 3.1 Definitions and problem statement
$G=(V^{l},E^{l})$ Road network, where $E^{l}$ represents the set of directed edges.
$V^{o},\,V^{p},\,V^{l}$ Sets of OD nodes, paths, and road segments, respectively.
$A^{o},\,A^{p},\,A^{l}$ Attribute matrices for OD nodes, paths, and road segments, respectively.
$U^{o},\,U^{p},\,U^{l}$ Sets of attribute nodes for OD nodes, paths, and road segments, respectively.
$\langle v_{c},v_{t},f\rangle$ OD flow: $f$ vehicles traveling from origin $v_{c}$ to destination $v_{t}$ within a time period.
$p:\langle v^{l}_{1},v^{l}_{2},\dots,v^{l}_{|p|}\rangle$ A path composed of $|p|$ road segments $v^{l}_{1},v^{l}_{2},\dots,v^{l}_{|p|}$.
$TG=(V,A,M)$ Trip Graph, where $V=V^{o}\cup V^{p}\cup V^{l}$, $A=A^{o}\cup A^{p}\cup A^{l}$, and $M=M_{o,p}\cup M_{p,l}\cup M_{l,l}$.
$M_{o,p}$ OD–Path traffic proportion matrix, with entries $m_{o,p}\in[0,1]$.
$M_{p,l}$ Path–Segment incidence matrix, where $m_{p,l}=\mathbb{I}\!\left(v^{l}\in p\right)$.
$M_{l,l}$ Topologically weighted segment–segment connectivity matrix.
$AG_{\phi}\!\left(V^{\phi}\!\cup\!U^{\phi},\,A^{\phi}\right)$ An attribute-guided graph for a specific node type $\phi$, where $\phi\in\{o,\,p,\,l\}$ corresponds to OD nodes, paths, and road segments, respectively.
$P_{i,j}^{TG}$ The transition probability from node $v^{\phi}_{i}$ to node $v^{\phi}_{j}$ on the TG, where $v^{\phi}_{i},\,v^{\phi}_{j}\in V$, and $\phi\in\{o,\,p,\,l\}$.
$P^{AG_{\phi}}_{i,k}$ The transition probability from node $v^{\phi}_{i}\in V^{\phi}$ to attribute $v^{\phi}_{k}\in U^{\phi}$ on the $AG_{\phi}$, where $\phi\in\{o,\,p,\,l\}$.
$Y=[y_{1},y_{2},\dots,y_{k}]$ Ground-truth node ranking list, where $k$ is the list length.
$\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{k}]$ Predicted node ranking list, where $k$ is the list length.
$\alpha\in[0,1]$ Walk bias factor in the TG.
$\beta\in[0,1]$ Probability of selecting a depth-first exploration strategy at each step of the random walk.
numnum, lenlen Parameters for generating numnum random walk sequences, each of length lenlen.
$h_{i}$ Final representation vector for a road segment.
Table 1 provides a summary of key notation and their definitions. This paper proposes a joint random walk-based representation learning framework on the TG and AG to improve the performance of the road segment ranking. This is achieved by integrating OD-driven functional structure and attribute-based functional similarity into the learned representations, which enables our method to more accurately capture long-range spatial dependencies in the traffic network.
@@ 3.2 Research objectives
(RO1) To design a joint heterogeneous graph random walk mechanism that generates context-rich sequences by sampling from both OD-driven functional structures and attribute-based similarities.
(RO2) To formulate a learning-to-rank framework that leverages the functional structure of road networks to achieve accurate prioritization of the most critical segments.
(RO3) To investigate whether the proposed model’s performance on road-network ranking tasks—relative to topology-based baselines—stems from its capacity to capture the underlying functional structure.
@@ 4 Methodology
This paper proposes HetGL2R, a deep learning framework (Figure 2) based on random walks on heterogeneous graphs, to improve traffic network ranking by integrating their functional structures with physical topology. The framework first extracts information from vehicle trajectory data to construct an OD-path-segment tripartite graph, which integrates the functional and topological structures of the road network into a unified representation. Simultaneously, a set of three attribute-guided bipartite graphs—one each for OD pairs, paths, and segments—are constructed to model functional similarities among nodes. Subsequently, the heterogeneous graph joint random walk algorithm, HetGWalk, is applied to these graphs to sample context-rich node sequences. These sequences are encoded using a Transformer encoder to learn road segment embeddings that capture long-range dependencies. Finally, a ranking learning module comprehensively evaluates and ranks the importance of road segments.
Figure 2: Framework of the proposed method.
@@ 4.1 HetGWalk algorithm
4.1.1 HetGWalk Overview
This subsection introduces the joint random walk algorithm, HetGWalk. The algorithm generates node sequences through probabilistic sampling on both the Trip Graph (TG) and the Attribute-guided Graph (AG): walks on the TG utilize OD-flow-path-segment connections to capture functional relationships and local topology of traffic networks, while walks on the AG leverage node-attribute relationships to encode functional similarities. The AG is introduced to enrich node context based on attributes, which increases sequence diversity and reduces the walks’ tendency to over-sample high-centrality nodes. By integrating these two types of walks, HetGWalk produces heterogeneous sequences that integrate functional and topological information, providing rich context for downstream representation learning. Figure 3 illustrates the sampling process.
The walk starts from the OD node and traverses all heterogeneous nodes in TG and all three AGs. To simultaneously capture both the local and global structures of TG, HetGWalk integrates depth-first exploration and breadth-first exploration. Given the current node $v^{\phi}_{i}$, we flip a coin with probability $\alpha$ of landing heads. If it lands on heads, we step in TG, with the probability of transitioning to $v^{\phi}_{j}$ given by:
@@ 4.1.1 HetGWalk Overview
where $\phi\in\{o,\,p,\,l\}$ denotes the node type, corresponding to OD nodes, paths, and road segments, respectively; $\beta$ is the probability of choosing depth-first exploration at each step of the walk; $i^{-}$ is the index of the predecessor node of $v^{\phi}_{i}$; $\tilde{P}_{i,j}$ represents the transition probability from $v^{\phi}_{i}\in V$ to $v^{\phi}_{j}\in V$ in a depth-first manner, which is calculated as follows:
$\tilde{P}_{i,j}^{TG}=\left[\begin{array}[]{ccc}0&M_{op}&0\\ 0&0&\bar{M}_{pl}\\ 0&(1-\varepsilon)\bar{M}_{pl}^{T}&\varepsilon M_{ll}\end{array}\right]\in\mathbb{R}^{|V|\times|V|}$ (3)
where, $\varepsilon$ denotes the probability of selecting segment nodes for the next step from the current segment node. If the coin lands tails up, the walk proceeds on the AG corresponding to the current node type in two steps:
The walk first transitions from the current entity node $v^{\phi}_{i}$ to one of its associated attribute nodes $u^{\phi}_{k}$. The probability of this transition, $P_{i,k}^{AG_{\phi}}$, is proportional to the importance of attribute $u^{\phi}_{k}$ to node $v^{\phi}_{i}$, specifically given by the normalized attribute value $\bar{a}^{\phi}_{k,i}$:
Figure 3: Joint random walk sampling procedure.
@@
This means that attributes contributing more to node $v^{\phi}_{i}$ have a higher probability of being sampled.
Next, the walk transitions from this attribute node $u^{\phi}_{k}$ to a new entity node $v^{\phi}_{j}$. The probability of this transition, $P_{k,j}^{AG_{\phi}}$, is determined by the attribute-value similarity between the target node $v^{\phi}_{j}$ and the previous entity node $v^{\phi}_{i}$ regarding attribute $u^{\phi}_{k}$. We employ a similarity measure based on the absolute difference:
@@
Input: TG, AG, $\alpha$, $f$, num, len, $\beta$
Output: The sampling sequence collections of $|V^{o}|$ nodes
2
Data: L1 norm to row-normalize $M^{T}$ and $A^{T}$ /* walk starts from an OD node */
3
4foreach $v^{o}_{t}\in V^{o}$ do
5 $S_{t}=\left\{\right\}$;
6 for $n=1$ to num do
7    for $l=1$ to len do
8       $s=\{v^{(t)}\}$;
9       $i\leftarrow t$;
10         flip a coin with probability $\alpha$ of landing heads;
11       if heads then
12            With probability $P_{i,j}^{TG}$, move to node $v^{(j)}$ in TG;
13            Append $v^{(j)}$ to $s$;
14
@@ 13            Append $v^{(j)}$ to $s$;
16            With probability $P^{AG_{\phi}}_{i,k}$, move to $u^{\phi}_{k}\in U^{\phi}$ in $AG_{\phi}$;
@@ 17            Append $u^{\phi}_{k}$ to $s$;
18            With probability $P^{AG_{\phi}}_{k,j}$, move to $v^{(j)}\in V^{\phi}$ in $AG_{\phi}$;
19            Append $v^{(j)}$ to $s$;
20
21       $i\leftarrow j$;
@@ 23     Append $s$ to $S_{t}$;
This design encourages the walk to jump to nodes that are similar to the origin node $v^{\phi}_{i}$ in terms of the specific attribute $u^{\phi}_{k}$, thereby capturing functional similarity even if the two nodes are not directly connected in the original Trip Graph.
Through this mechanism, the attribute-guided walk incorporates both the influence and similarity of node attributes into the transition process. The overall probability of reaching $v^{\phi}_{j}$ from $v^{\phi}_{i}$ is computed as a weighted aggregation over all possible paths passing through attribute nodes $u^{\phi}_{k}$, effectively constructing an implicit similarity network enriched with attribute semantics. The implementation details of HetGWalk are described in Algorithm 1.
Line 1: The algorithm iterates over each OD node $v^{o}_{t}$ in the set $V^{o}$, which serves as the starting point for random walks. Line 2: Initializes an empty collection $S_{t}$ to store the generated sequences for the current OD node $v^{o}_{t}$. Line 3: For each starting node, the algorithm generates numnum independent random walk sequences. Line 4: Each sequence grows incrementally until reaching the predefined length lenlen. Line 7: At each step, a random decision is made: with probability $\alpha$, the walk continues in the Trip Graph; otherwise, it proceeds to the Attribute-guided Graph. Lines 9-10: If the decision selects TG traversal, the walk moves to a neighboring node $v^{(j)}$ according to the transition probability $P_{i,j}^{TG}$ (Eq. 2) and appends it to the sequence.
Lines 12-15: If AG traversal is selected, the walk executes a two-stage process: First transitions from the current node $v^{(t)}$ to an attribute node $u^{\phi}_{k}$ using probability $P_{i,k}^{AG_{\phi}}$ (Eq. 4). Then immediately moves to entity node $v^{(j)}$ using probability $P_{k,j}^{AG_{\phi}}$ (Eq. 5). Both the attribute node and the new entity node are appended to the sequence. Line 16 Updates the current node position to continue the walk from $v^{(j)}$. Line 17: Stores the completed sequence $s$ in the collection $S_{t}$. Line 18: Returns all generated sequences for all starting OD nodes.
@@ 4.1.2 Theoretical Properties of HetGWalk
To rigorously analyze the behavior of our proposed heterogeneous graph representation learning framework, we formalize the properties of attribute-guided graphs (AGs) and joint random walks on the trip graph (TG) and AGs.
Definition 7 (Type-Specific Attribute Propagation Operators): For a node $v_{i}^{\phi}\in V^{\phi}$, $\phi\in\left\{o,p,l\right\}$, the propagation operator $\rho_{i}^{\phi}(j)$ measures the probability of influence propagation from $v_{i}^{\phi}$ to another node $v_{j}^{\phi}$ via attributes:
$\color[rgb]{0,0,0}\rho_{i}^{\phi}(j)=\sum_{k\in U^{o}}\bar{a}_{k,i}^{\phi}\cdot\left(1-\left|\bar{a}_{k,j}^{\phi}-\bar{a}_{k,i}^{\phi}\right|\right),\quad v_{j}^{\phi}\in V^{\phi}$} (6)
@@
Theorem 1 (Type-Aware Dynamic Semantic Propagation): For any node type $\phi\in\left\{o,p,l\right\}$, attribute-guided random walk transitions on $AG_{\phi}$ implement a type-specific semantic propagation process. Specifically, when the walk is currently at node $v_{i}^{\phi}$, the total transition probability from $v_{i}^{\phi}$ to any node $v_{j}^{\phi}$ is exactly equal to the propagation operator $\rho_{i}^{\phi}$.
@@
$\color[rgb]{0,0,0}P_{i,j}^{AG_{\phi}}=\sum_{k\in U^{\phi}}P_{i,k,j}^{AG_{\phi}}=\sum_{k\in U^{\phi}}\bar{a}_{k,i}^{\phi}\cdot\left(1-\left|\bar{a}_{k,j}^{\phi}-\bar{a}_{k,i}^{\phi}\right|\right)=\rho_{i}^{\phi}(j)$} (8)
@@
Discussion: The two theorems collectively demonstrate that TG and AGs provide complementary guidance for heterogeneous random walks. Specifically, the multiple AGs enable type-specific modeling of functional similarities, allowing transition probabilities to capture distinct functional semantics—such as OD-flow intensity, path-level relationships, and segment attributes. Meanwhile, the TG preserves multi-hop, cross-type functional dependencies, ensuring that walks capture essential functional-structural relationships among OD pairs, paths, and segments. Together, the TG and AGs form a coordinated mechanism that biases walks toward nodes that are both semantically relevant and structurally dependent. As a result, the generated heterogeneous sequences encode rich functional and structural contexts, providing a theoretical basis for learning embeddings that integrate both functional and structural dependencies.

### v3

@@ 3.1 Definitions and problem statement
G,TG,AGG,\,TG,\,AG road network graph, trip graph, attribute-guide graph
o,p,lo,\,p,\,l OD type, path type, segment type
$\phi$ node type, $\phi\in\{o,\,p,\,l\}$
$V^{\phi},\,v^{\phi}$ set of nodes of type $\phi$, a node of type $\phi$
$A^{\phi}$ attribute matrix for nodes of type $\phi$
$U^{\phi},\,u^{\phi}$ set of attribute nodes of type $\phi$, an attribute node of type $\phi$
V,A,UV,\,A,\,U $V^{o}\cup V^{p}\cup V^{l}$, $A^{o}\cup A^{p}\cup A^{l}$, $U^{o}\cup U^{p}\cup U^{l}$
$\rho$ a path
$\mathcal{P}_{i}$ set of paths containing segment $i$
$a_{ij}^{\phi},\,\bar{a}_{ij}^{\phi}$ $j$-th attribute value of node $i$ of type $\phi$ and its normalized form
$M^{o,p}$ OD–Path association matrix
$M^{p,l},\,{\color[rgb]{0,0,0}\bar{M}^{p,l},\,\bar{M}^{\prime p,l}}$ path–segment incidence matrix and its row- and column-normalized forms
$M^{l,l},{\color[rgb]{0,0,0}\bar{M}^{l,l}}$ segment–segment connectivity matrix and its row-normalized form
$H$ list of road segment representations used for ranking
$K$ number of segments in $H$
$h_{i}$ representation of segment $i$
num,lennum,\,len number of generated sequences per OD, length of each sequence
$\alpha$ probability of choosing TG-based transitions in the walk
$\beta$ trade-off parameter between depth-first and breadth-first transitions on the TG
$\varepsilon$ trade-off parameter between segment–segment and path–segment transitions on the TG
$Y,\,\hat{Y}$ ground-truth and predicted rank position vectors of a segment list
$y_{i},\,\hat{y}_{i}$ ground-truth and predicted rank positions of segment $i$
In summary, the above definitions provide the structural and semantic foundations for modeling road networks from a functional perspective. By jointly considering OD flows, paths, and attribute-based similarities, they enable representation learning methods to go beyond purely topological dependencies and capture the functional interactions that govern long-range spatial influence in traffic networks. Table 1 provides a summary of key notation and their definitions.
@@ 3.2 Research objectives
Building upon the problem definition in Section 3.1, this study is guided by the following research objectives:
(RO1) To design a joint heterogeneous graph random walk mechanism that generates context-rich sequences by jointly sampling OD-induced functional structures and attribute-based similarities, enabling the model to encode both functional interactions and semantic consistency among road segments.
(RO2) To formulate a learning-to-rank framework that leverages the functional structure of road networks-induced by OD flows and path interactions-to achieve accurate prioritization of the critical road segments.
(RO3) To examine the contribution of functional-structure modeling to road-network ranking performance through comparisons with topology-driven baselines, thereby clarifying the role of OD-induced functional dependencies beyond purely structural cues.
These objectives directly inform the design of the proposed HetGL2R framework and the subsequent experimental evaluation presented in the following sections.
@@ 4 Methodology
Figure 2: Framework of the proposed HetGL2R method. The framework consists of four main stages: heterogeneous graph construction, sequence generation via HetGWalk, feature encoding, and learning to rank. Yellow, green, and blue nodes represent origin–destination (OD) pairs, paths, and road segments, respectively. Solid arrows indicate the main data flow, while dashed arrows denote the saving and updating processes during model training. Specifically, HetGWalk samples from the trip graph and attribute-guided graphs to generate heterogeneous sequences, which are encoded by a Transformer-based encoder and subsequently used for listwise ranking.
This paper proposes HetGL2R, a deep learning framework (Figure 2) based on random walks on heterogeneous graphs, to improve traffic network ranking by integrating their functional structures with physical topology. The framework first extracts information from vehicle trajectory data to construct an OD-path-segment tripartite graph, which integrates the functional and physical structures of the road network into a unified representation. Simultaneously, a set of three attribute-guided bipartite graphs—one each for OD pairs, paths, and segments—are constructed to model functional similarities among nodes. Subsequently, the heterogeneous graph joint random walk algorithm, HetGWalk, is applied to these graphs to sample context-rich node sequences. These sequences are encoded using a Transformer encoder to learn road segment embeddings that capture long-range dependencies. Finally, a ranking learning module comprehensively evaluates and ranks the importance of road segments.
@@ 4.1 HetGWalk algorithm
4.1.1 Joint Random Walk Procedure
This subsection introduces HetGWalk, a joint random walk algorithm designed to generate heterogeneous node sequences by sampling over both the Trip Graph (TG) and the Attribute-guided Graphs (AGs). Walks on the TG exploit OD–path–segment relationships to capture OD-induced functional dependencies together with the underlying network topology, while walks on the AGs leverage node–attribute associations to encode functional similarities among nodes of the same type. By enriching node context with attribute information, the AGs increase sequence diversity and alleviate the tendency of random walks to oversample high-centrality nodes. By integrating these two complementary sampling mechanisms, HetGWalk produces context-rich heterogeneous sequences that jointly encode functional and physical structures, providing informative inputs for downstream representation learning. Figure 3 illustrates the overall sampling procedure.
Figure 3: Joint random walk sampling procedure. Joint sampling is performed over the Trip Graph (TG) and multiple Attribute-Guided Graphs (AGs). The mixing coefficient $\alpha$ controls the probability of switching between TG-based and attribute-guided transitions at each step. Red arrows show an example sampling trajectory. On the right, numnum denotes the number of sequences sampled per node, and lenlen indicates the fixed sequence length.
Each walk starts from an OD node and probabilistically traverses heterogeneous nodes in the TG as well as the three AGs corresponding to OD pairs, paths, and road segments. To jointly exploit the local and global functional structures encoded in the TG, HetGWalk integrates depth-first and breadth-first exploration strategies to balance OD-consistent path continuity and cross-path, cross-OD exploration. At each step, given the current node $v^{\phi}_{i}$, where $\phi\in\left\{o,p,l\right\}$ denotes the node type (OD pair, path, or road segment), a Bernoulli trial with probability $\alpha$ determines whether the next transition is sampled from the TG or from the AGs. When sampling on the TG, the transition probability is defined as
@@ 4.1.1 Joint Random Walk Procedure
where $\beta$ controls the trade-off between depth-first and breadth-first exploration, $i^{-}$ denotes the predecessor of node $v^{\phi}_{i}$ in the walk, and $\tilde{P}_{i,j}^{TG}$ represents the depth-first transition probability on the TG. This probability depends on the types of the current and next nodes and is calculated as
$\tilde{P}_{i,j}^{TG}=\left[\begin{array}[]{ccc}0&M^{o,p}&0\\ 0&0&\bar{M}^{p,l}\\ 0&(1-\varepsilon)\bar{M}^{\prime p,l}&\varepsilon\bar{M}^{l,l}\end{array}\right]\in\mathbb{R}^{|V|\times|V|}$ (3)
where $\bar{M}^{p,l}$ and $\bar{M}^{\prime p,l}$ are the row- and column-normalized path–segment incidence matrices, respectively, and $\bar{M}^{l,l}$ is the row-normalized segment–segment connectivity matrix defined in Definition 4. The parameter $\varepsilon\in[0,1]$ biases the walk toward segment–segment transitions along OD-induced downstream dependencies, while $1-\varepsilon$ encourages returning to path nodes to preserve OD–path–segment consistency.
When the walk transitions to an AG, it follows an entity–attribute–entity pattern. First, the walk moves from the current entity node $v^{\phi}_{i}$ to one of its associated attribute nodes $u^{\phi}_{k}$. The transition probability $P_{i,k}^{AG_{\phi}}$ is proportional to the importance of attribute $u^{\phi}_{k}$ to node $v^{\phi}_{i}$, and is defined as the normalized attribute value $\bar{a}^{\phi}_{k,i}$:
@@
This design ensures that attributes contributing more significantly to node $v^{\phi}_{i}$ are sampled with higher probability. Next, the walk transitions from the selected attribute node $u^{\phi}_{k}$ to a new entity node $v^{\phi}_{j}$. The corresponding transition probability $P_{k,j}^{AG_{\phi}}$ is determined by the functional similarity between nodes $v^{\phi}_{i}$ and $v^{\phi}_{j}$ with respect to attribute $u^{\phi}_{k}$, defined as:
@@
Input: TG, $AG_{o},AG_{p},AG_{l}$, walk length lenlen, walks per OD numnum
Output: The sampling sequence collections of $|V^{o}|$ nodes
@@
[…]
3 for each OD node $v^{o}_{i}\in V^{o}$ do
4 for $n=1$ to numnum do
5    $s\leftarrow[v^{o}_{i}]$, predecessor $v^{-}\leftarrow\text{null}$ ;
6    for $t=2$ to lenlen do
7         current node $v^{\phi}_{i}\leftarrow s[t-1]$ with type $\phi$ ;
8         Sample $z\sim\text{Bernoulli}(\alpha)$ ;
9       if $z=1$ then
10            Sample $v^{\phi}_{j}$ from TG using $P_{i,j}^{TG}$ (Eq. (3)) given ($v^{\phi}_{i},v^{-}$) ;
11
@@ 8         Sample $z\sim\text{Bernoulli}(\alpha)$ ;
13            Sample attribute $u^{\phi}_{k}$ using $P^{AG_{\phi}}_{i,k}$ (Eq. (4)) ;
@@ 14            Append $u^{\phi}_{k}$ to $s$ ;
15            Sample $v^{\phi}_{j}$ using $P^{AG_{\phi}}_{k,j}$ (Eq. (5)) ;
16            Append $v^{\phi}_{j}$ to $s$ ;
17
18       $v^{-}\leftarrow v^{\phi}_{i}$ ;
19
20     Add $s$ to $S$ ;
21
@@ 20     Add $s$ to $S$ ;
The resulting similarity scores are normalized over candidate entity nodes to form a valid transition distribution. This formulation encourages the walk to transition to entity nodes that are functionally similar to the origin node $v^{\phi}_{i}$ under the selected attribute, even in the absence of direct connectivity in the original TG.
Overall, the probability of reaching node $v^{\phi}_{j}$ from $v^{\phi}_{i}$ via the AG is computed as a weighted aggregation over all possible attribute-mediated paths. Through this mechanism, the walk implicitly induces an attribute-aware similarity structure enriched with functional semantic, enabling the model to capture functional relationships beyond explicit graph connectivity. The implementation details of HetGWalk are described in Algorithm 1.
Lines 1–2 initialize an empty sequence set and iterate over all OD nodes. Lines 3–4 generate numnum walk sequences for each OD node and initialize each walk with the OD node as the starting point. Lines 5–6 iteratively extend the walk by identifying the current node at each step. Line 7 samples a Bernoulli random variable to decide whether the next transition is taken on the TG or on an AG. Lines 8–9 sample the next node from the TG according to Eq. (2). Lines 10–13 sample the next node via the corresponding AG using an entity–attribute–entity transition. Lines 14–15 append the sampled node(s) to the current walk and update the predecessor. Lines 16 add the completed walk to the sequence set, and Line 17 outputs the set of generated heterogeneous walk sequences.
@@ 4.1.2 Theoretical Properties of HetGWalk
In this subsection, we formalize the properties of AGs and joint random walks on the TG and AGs for the proposed heterogeneous graph representation learning framework.
Definition 7 (Type-Specific Attribute Propagation Operators): For a node $v_{i}^{\phi}\in V^{\phi}$, $\phi\in\left\{o,p,l\right\}$, the propagation operator $\vartheta_{i}^{\phi}(j)$ measures the probability of influence propagation from $v_{i}^{\phi}$ to another node $v_{j}^{\phi}$ via attributes:
$\color[rgb]{0,0,0}\vartheta_{i}^{\phi}(j)=\sum_{k\in U^{o}}\bar{a}_{k,i}^{\phi}\cdot\left(1-\left|\bar{a}_{k,j}^{\phi}-\bar{a}_{k,i}^{\phi}\right|\right),\quad v_{j}^{\phi}\in V^{\phi}$} (6)
@@
Theorem 1 (Type-Aware Dynamic Semantic Propagation): For any node type $\phi\in\left\{o,p,l\right\}$, attribute-guided random walk transitions on $AG_{\phi}$ implement a type-specific semantic propagation process. Specifically, when the walk is currently at node $v_{i}^{\phi}$, the total transition probability from $v_{i}^{\phi}$ to any node $v_{j}^{\phi}$ is exactly equal to the propagation operator $\vartheta_{i}^{\phi}$.
@@
$\color[rgb]{0,0,0}P_{i,j}^{AG_{\phi}}=\sum_{k\in U^{\phi}}P_{i,k,j}^{AG_{\phi}}=\sum_{k\in U^{\phi}}\bar{a}_{k,i}^{\phi}\cdot\left(1-\left|\bar{a}_{k,j}^{\phi}-\bar{a}_{k,i}^{\phi}\right|\right)=\vartheta_{i}^{\phi}(j)$} (8)
@@
Discussion: The two theorems collectively demonstrate that TG and AGs provide complementary guidance for heterogeneous random walks. Specifically, the multiple AGs enable type-specific modeling of functional similarities, allowing transition probabilities to capture distinct functional semantics—such as OD-flow intensity, path-level relationships, and segment attributes. Meanwhile, the TG preserves multi-hop, cross-type functional dependencies, ensuring that walks capture essential functional-structural relationships among OD pairs, paths, and segments. Together, the TG and AGs form a coordinated mechanism that biases walks toward nodes that are both semantically relevant and structurally dependent. This coordination directly affects the distribution and context of sampled nodes, increasing the exposure of functionally critical segments during training. As a result, segments that are important under specific OD flows or path interactions are embedded in richer and more informative contexts, yielding representations that are better suited for downstream ranking.